OpenStudy (anonymous):

Simpilfy 1. 4x/9 + x-10/9 Show work.

OpenStudy (anonymous):

You need to gather the like terms: there is only one constant term, but there are two terms that involve x: 4x/9 and x \[\frac{4x}{9}=\frac{4}{9}x\] so we have 4/9 + 1 = 1 and 4/9 x

OpenStudy (anonymous):

so the answer os 4/9 x?

OpenStudy (anonymous):

no: \[\frac{4}{9}x+x-\frac{10}{9}=1\frac{4}{9}x-\frac{10}{9}\]